### v1

# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities

Abstract
In recent years, extensive research has been conducted to explore the utilization of machine learning algorithms in various direct-detected and self-coherent short-reach communication applications. These applications encompass a wide range of tasks, including bandwidth request prediction, signal quality monitoring, fault detection, traffic prediction, and digital signal processing (DSP)-based equalization. As a versatile approach, machine learning demonstrates the ability to address stochastic phenomena in optical systems networks where deterministic methods may fall short. However, when it comes to DSP equalization algorithms like feed-forward/decision feedback equalizers (FFE/DFE) and Volterra-based nonlinear equalizers, their performance improvements are often marginal, and their complexity is prohibitively high, especially in cost-sensitive short-reach communications scenarios such as passive optical networks (PONs). Time-series machine learning models offer distinct advantages over frequency domain models in specific contexts. They excel in capturing temporal dependencies, handling irregular or nonlinear patterns effectively, and accommodating variable time intervals. Within this extensive survey, we outline the application of machine learning techniques in short-reach communications, specifically emphasizing their utilization in high-bandwidth demanding PONs. Notably, we introduce a novel taxonomy for time-series methods employed in machine learning signal processing, providing a structured classification framework. Our taxonomy categorizes current time series methods into four distinct groups: traditional methods, Fourier convolution-based methods, transformer-based models, and time-series convolutional networks. Finally, we highlight prospective research directions within this rapidly evolving field and outline specific solutions to mitigate the complexity associated with hardware implementations. We aim to pave the way for more practical and efficient deployment of machine learning approaches in short-reach optical communication systems by addressing complexity concerns.
keywords:
machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities
[…]
1 Introduction
Short-reach optical transmission systems have gained substantial attraction owing to their remarkable attributes of high bandwidth and low latency Durisi (2016). In the evolving landscape of communication technologies, short-reach optical communication has emerged as an essential domain, driven by the increasing demand for high-speed data transfer in applications such as inter-data centers Kapoor (2012), access/local area networks and industrial automation XieY (2022). This increasing demand requires efficient, low-latency communication systems tailored to short-reach scenarios, typically up to 100 km. While long-haul, optical communication has been immersive in data transmission, but its applicability encounters challenges when adapting to the constraints of shorter distances. This is mainly due to physical and technical limitations that prevents its seamless integration into contemporary networking environments characterized by the need for rapid, energy-efficient, and cost-effective data transfer over limited distances. Passive optical networks (PONs) have emerged as a cost-efficient solution for access networks, effectively addressing challenges and meeting the requirements of modern networking scenarios. PONs optimize data transfer efficiency within constrained distances, providing a streamlined alternative to point-to-point communications.
Despite their efficiency, the capacity of PONs and short-reach optical systems, in general, is limited by intrinsic properties of optical components, as well as the fiber itself, for distortions such as chromatic dispersion (CD), nonlinearity, and polarization mode dispersion (PMD) Wu, Q. (2023). For instance, intensity-modulation and direct-detection (IMDD) performance is hindered by power fading attributed to fiber CD Ranzini (2019).
Ongoing research endeavors are dedicated to advancing optical detection schemes to overcome these limitations and facilitate high speed in both short-reach and long-haul optical communication networks D. Che (2016). For instance, the regeneration of coherent optical systems in the last decade has marked a breakthrough by transcending intensity-only modulation Li (2022). Coherent systems employ external modulators to imprint complex baseband signals to the optical field. The optical coherent receiver, equipped with phase diversity, linearly recovers signals and compensates for fiber impairments through digital signal processing (DSP) S. J. Savory (2007). Coherent technology enables the transmission of advanced modulation formats and polarization multiplexing to increase the signal bit rate significantly. Additionally, it triggers dense wavelength division multiplexing (DWDM) and super-channels, propelling long-haul optical networks into a capacity realm beyond multi-Terabit/s. C. R. S. Fludger (2008).
Except for traditional homodyne-coherent technology, coherent communication strategies include diverse techniques, such as phase detection through heterodyne detection. While this approach has its merits Delange (1968), a notably favored incoherent approach such as IMDD is practically preferred due to its inherent simplicity and cost-effectiveness in short-reach communications Khan (1997); Huang (2022).
In contrast to coherent transmission, IMDD operates by encoding information into the intensity of the optical signal, with the modulation signal being real-valued and positive Huang (2022). The implementation of IMDD eliminates the need for complex optical components and local oscillators, reducing hardware complexity. Additionally, IMDD systems are less susceptible to phase noise and polarization-related issues, making them robust and practical for scenarios where cost efficiency and simplicity are paramount. The advantages of IMDD, including its ease of implementation and resilience, position it as a compelling choice for optimizing short-distance optical communication Huang (2022). In addition, operational and safety factors can impose restrictions on the peak and average values of the modulated signal in IMDD. These constraints introduce distinct characteristics to its operational framework Kartal (2011). Various models, such as the Poisson channel, square-root Gaussian channel, and Gaussian channel with input-dependent noise, among others, exist to rapidly assess and characterize IMDD systems Tsiatmas (2013); Moser (2012); Safari (2015). Each of these models provides valuable insights into the intricate dynamics of this communication paradigm, enabling quick and efficient analysis. Recently, in short-reach communication systems, there has been a shift towards replacing the traditional DSP functionality with machine learning (ML) as the sole processing mechanism. In contrast to conventional methodologies that heavily depend on intricate analog components Fadlullah (2011), this strategy seamlessly integrates ML within the DSP framework Yi et al. (2018), providing a flexible and adaptable solution for enhancing optical communication over short distances. It has been reported in Kaur (2021) that the combination of ML with DSP techniques allows these systems to dynamically adapt and optimize signal parameters, effectively addressing impairments and variations in real time without requiring intricate hardware adjustments.
Furthermore, the integration of ML exclusively within the DSP domain brings streamlined and efficient implementation processes to short-reach communication systems. This approach represents a paradigm shift as it not only reduces costs associated with complex hardware setups but also emphasizes the effectiveness of intelligent signal processing solutions in short-reach optical transmission Simeone (2018); Yi et al. (2018); Rodrigues (2019).
In this survey, we examine the significant advancements in short-reach communication research over the last decade. Firstly, we summarize several key research tasks (Section 2) and then focus on the equalization problem, introducing benchmark DSP methods (Section 3) and ML algorithms (Section 4). Subsequently, we categorize recent sequence models in the ML field (Section 5), dividing them into convolution-based models, transformer-based models, and Fourier-based neural networks. We explore each method’s advantages, disadvantages, and complexities in addressing the equalization problem. In the last section, we present an overview of the model compression field, outlining two approaches to compressing models. We view these approaches as potential solutions for tackling hardware complexity concerns.
Our primary contribution lies in summarizing the past research on ML implementation in short-reach communications for various applications. Specifically, our contributions are the following:
  1. 1.
    We review existing deep learning models, providing a comprehensive understanding of their principles, characteristics, and hypothesis classes. This facilitates an in-depth exploration for researchers seeking supervised neural-network-based ML models suitable for their specific applications.
  2. 2.
    We highlight the features and complexities of these models, elucidating recent developments in the field of deep learning. This information is valuable for researchers interested in delving deeper into research and staying abreast of current advancements.
  3. 3.
    We discuss the current limitations and research gaps in the ongoing development of deep learning, addressing the challenges posed by these factors in real-world applications. Moreover, we proffered constructive insights regarding model selection and prospective trajectories.
  4. 4.
    Given the prevailing backdrop of high hardware complexity, we introduce a solution in the field of deep learning: the principles of model compression. We present existing works that employ this approach within the optical communication field, aiming to inspire more researchers to pursue research in this domain.
2 Applications in Short-Reach Systems
After systematically organizing recent literature in the past few years, we categorize machine learning-based short-reach system research into four classes based on application tasks: Bandwidth Request and Prediction, Subcarrier Allocation, Equalization and Fault Detection. We clarify the physical and mathematical aspects of their respective tasks, enumerate several recent works, and provide a summary of current advancements.
Bandwidth Request and Prediction: It aims to leverage network information to predict future bandwidth availability and enable its utilization by related applications. In mathematical terms, the real-time bandwidth forecast at a specific time (t) involves estimating the available bandwidth that will be accessible in the immediate future (t + $\tau$) Qi-yuZ (2012). One proposed method, known as predictive-dynamic bandwidth allocation (P-DBA), utilizes this concept to predict high-priority traffic during waiting periods, resulting in reduced latency and packet loss rates within Gigabit PON (GPON) Qi-yuZ (2012). Another approach demonstrated in PanagiotisS (2016) leverages the k-nearest neighbor algorithm to predict additional bandwidth requirements for each optical network unit (ONU) in a PON. This adaptive learning-based approach dynamically adjusts the k value based on real-time traffic conditions, showcasing the adaptability of ML in optimizing bandwidth allocation PanagiotisS (2016). Artificial neural networks (ANNs) have also shown promise in achieving flexible bandwidth allocations across various application scenarios, particularly emphasizing low-latency objectives LihuaR (2018); YiL (2019). For example, feed-forward-based ANNs, explored in Mikaeil (2018), are utilized to predict packet arrivals in time-division multiple access (TDMA) ONUs, effectively reducing additional DBA processing delays Mikaeil (2018). Furthermore, in Ye (2018), XgboostXgboost (2016) is employed to predict bandwidth requests for ONUs in Ethernet PON (EPON), optimizing bandwidth utilization across polling periods. This study introduces a dynamic wavelength and bandwidth assignment scheme for time and WDM (TWDM) PONs, incorporating regression techniques for efficient resource allocation Ye (2018). Collectively, these studies highlight the versatility of ML approaches in addressing bandwidth prediction and request challenges, paving the way for more adaptive and efficient short-reach optical communication systems Qi-yuZ (2012); PanagiotisS (2016); LihuaR (2018); YiL (2019); Mikaeil (2018); Ye (2018).
Subcarrier Allocation: The optimization of bandwidth allocation for enhanced spectral efficiency has led to increased interest in subcarrier allocation for PONs. This approach involves mathematically formulating the allocation problem as an integer linear programming (ILP) task, which includes tasks such as optimizing wavelength configurations, assigning subcarriers to transmitters, and minimizing lost traffic and energy costs. To address this challenge, deep reinforcement learning has emerged as a promising technique that enables dynamic subcarrier sharing among ONUs, facilitating efficient DBA. At the medium access control (MAC) layer, the dynamic subcarrier allocation (DSA) algorithm schedules ONU upstream transmissions by considering instantaneous bandwidth requirements and existing traffic conditions KonstantinosK (2012) . This showcases the adaptability of ML in resource scheduling. Several studies focus on algorithm-level cost reduction and two-dimensional resource scheduling for orthogonal frequency-division multiplexing (OFDM)-PONs including KonstantinosK (2012); WansuL (2017); MeihuaB (2013). These DSA algorithms address challenges related to latency, throughput, and energy efficiency, highlighting the versatility of ML in enhancing subcarrier allocation strategies ZhuM (2022). Moreover, the integration of traffic prediction technology and fair-aware DSA algorithms, as proposed in YumikoS (2017) and ZhuM (2022), further enhances the performance of subcarrier allocation in short-reach optical communication systems. These advancements improve the efficiency and adaptability of subcarrier allocation by applying ML methodologies YNakayama (2022).
Power Budget Problem: The electric power budgeting problem is focusing on predicting energy consumption for future time intervals with the given historical data on power consumption and relevant environmental factors such as weather conditions, user behavior, and equipment efficiency, the goal is to predict for future time intervals. However, the large-scale systematic ML development is largely constrained by the limited number of publicly available datasets, and recent work has provided a basic data construction process and baseline model Alberto C (2023).
Equalization: The objective of this task is to minimize fiber-induced distortions by employing post-processing techniques that compensate for linear effects, such as chromatic dispersion (CD). Mathematically, the equalizer optimizes the function $f(x)$ to ensure that the equalized output sequence $y$ closely approximates the input signal. Performance evaluation primarily relies on the bit error ratio (BER). In PON systems, the utilization of post-equalizers that employ shallow-based deep learning models shows promise in mitigating nonlinear distortions, especially in scenarios involving modulator nonlinearities or high launched optical power to meet stringent power budgets Li (2022). In the context of PONs, these situations can give rise to self-phase modulation (SPM) in single-channel scenarios or, in the case of multi-channel PONs, additional phenomena such as cross-phase modulation (XPM) and four-wave mixing (FWM). In IMDD PON, low-complexity ANN-based equalizers have demonstrated comparable performance to Volterra-based equalizers in PAM4 systems ZhiweiChen (2022). While post-equalization techniques have proven effective, the computational complexity at the receiver side poses challenges for the ONU. To address this issue, strategies for centralized pre-equalization at the transmitter side have been proposed. Examples of these strategies include memory polynomial-based pre-equalizers ZhiweiChen (2022) and trained NN-based pre-equalizers XueL (2021). These methods enhance equalization effectiveness while maintaining the simplicity of the DSP at the ONU receiver. These advancements highlight the potential of ML in optimizing equalization strategies in PONs and other short-reach systems. This survey provides a comprehensive overview of existing equalization techniques,focusing on deep learning methods that have demonstrated effectiveness but have not yet been widely adopted.
Fault Detection: Short-reach optical communication systems, including PONs, are susceptible to failures such as fiber cuts, equipment failures, power outages, and natural disasters and ONU transceiver malfunctions Abdelli (2023). The impact of service disruptions can result in significant financial losses for service providers. Identifying faulty ONUs presents challenges, especially in scenarios where nearly equidistant branch terminations lead to overlapping reflections, making it difficult to pinpoint the exact defective branch Abdelli (2023). Conventional monitoring approaches become less reliable as PON systems grow in complexity. Recent advancements in ML-enabled proactive fault monitoring offer promising solutions to ensure stable network operation. ML-based fault prediction algorithms utilize past network fault data to discover underlying patterns and similarities among events. By doing so, these algorithms enhance the detection of optical network problems and facilitate proactive repairs, thereby avoiding potential issues from occurring.. Several research papers propose using ML algorithms for monitoring management in optical networks. Notably, technologies like random forest and ANN algorithms have been employed to continuously monitor the BER, predict network component failures, and assess fault severity APVela (2018). Wang et al. introduced a hybrid approach combining double exponential smoothing and support vector machines for equipment failure prediction in software-defined metropolitan area networks Wang (201). Bayesian network-based models have also been developed for diagnosing PON faults Wang (201).
3 DSP for Signal Equalization in Communication Systems
In this section, we provide an overview of conventional signal equalization techniques, ranging from basic zero-forcing equalization to more advanced approaches such as feed-forward equalizers (FFE), decision-feedback equalizers (DFE), Viterbi and Volterra equalizers, as well as adaptive equalizers. We discuss the advantages and limitations of these techniques and present a comparative analysis alongside ML models. We briefly introduce the principles of these methods and provide a complexity analysis for each of them in combination with traditional sequential ML methods in the Table 1.
Zero Forcing (ZF): Zero forcing is an LE derived by minimizing inter-symbol interference (ISI). A groundbreaking study Tufts (1965) has established the analytical foundation for optimal zero-forcing and minimum mean squared error (MSE) equalization in channels with additive white noise and specified frequency response. The study further demonstrates that an optimal LE can be implemented as a cascade of filters, with taps spaced at symbol intervals. However, when the channel effect exhibits a deep frequency response “valleys,” equalization will yield poor performance due to the noise enhancement problem.
Feed Forward Equalizer: FFE RagaLasya (2017) mitigates ISI in communication channels by processing the received signal forwardly without feedback. Its simplicity makes it suitable for systems where feedback is challenging or undesired.
Decision Feedback Equalizers: Due to the noise enhancement, DFE is designed to reduce ISI by subtracting already-known symbols. In this way, ISI from already detected symbols is eliminated. Adaptation of the forward and feedback filters of DFEs follow the same pattern as for LEs WilliamsonD (1992). The disadvantage is that it could induce accumulated errors associated with feeding back filtered incorrect detector decisions.
Viterbi equalizer: The Viterbi equalizer seeks to estimate the most likely sequence of transmitted symbols, given the received sequence. By constructing a trellis diagram where nodes represent possible transmitted symbols and transitions denote potential channel transitions, the Viterbi algorithm dynamically optimizes path metrics to identify the most probable sequence. This process involves state transition probabilities and meticulous calculations to mitigate the impact of channel impairments. Mathematically, the Viterbi equalizer applies the Viterbi algorithm, which belongs to the dynamic programming algorithm for finding the most likely sequence of hidden states in a hidden Markov model. The time complexity of the Viterbi equalizer is determined by the Viterbi algorithm, which is influenced by the length of the input sequence and the number of states, making it $O(T\cdot N^{2})$, where $T$ denotes the length the sequence and $S$ refers to the number of the hidden states Forney (2005).
Volterra equalizer: Is a nonlinear equalizer used in optical communication systems to compensate for nonlinear distortions introduced by the fiber channel volterra (2103). In the context of PAM4 optical systems, a Volterra equalizer can be employed to mitigate the nonlinear effects that arise due to the intensity-dependent response of the optical components. The Volterra equalizer is based on the Volterra series expansion, which allows for the modeling of nonlinear systems. The key idea is to use a set of nonlinear filters, known as Volterra kernels, to capture the nonlinear characteristics of the channel. In the case of PAM4 optical systems, the Volterra equalizer is typically designed to handle the specific nonlinearities associated with the modulation format and the optical components employed. The structure of a Volterra equalizer consists of multiple stages, each representing a different order of nonlinearity. The first stage corresponds to the linear equalizer, which performs initial equalization to address linear distortions. Subsequent stages of the Volterra equalizer capture and compensate for higher-order nonlinear distortions. These stages involve nonlinear filters that take multiple past symbols as inputs and produce outputs based on their interaction. The number of stages and the complexity of the Volterra equalizer depend on the specific system requirements and the level of nonlinear distortions present. The coefficients of the Volterra kernels are typically adapted or optimized using algorithms such as the least mean squares (LMS) or recursive least squares (RLS) algorithms. These algorithms iteratively adjust the coefficients based on the error between the equalized signal and the desired signal, aiming to minimize the distortion and improve the overall system performance.
Adaptive filtering: Adaptive filtering Malik (2011) is used in communication systems where channel characteristics vary over time. The mathematical interpretation involves using an adaptive algorithm that iteratively modifies the filter parameters to minimize the error signal between the desired output and the actual output, enabling the filter to adapt to changing input conditions. The actual convergence time and the total time complexity over multiple iterations depend on the convergence behavior of the specific algorithm and its sensitivity to the input data. Assuming t taps, the total time complexity for updating all coefficients is O(t). FFE and DFE are regarded as adaptive filtering versions designed explicitly for short-reach communications.
Table 1: Complexity Analysis for DFE, FFE, LE, Adaptive Filtering and Viterbi algorithms. $t$ refers to the number of the taps. $N$ in Viterbi denotes the number of the hidden states.
Models DFE FFE LE Adaptive Filtering Viterbi
Train $O(t)$ $O(t)$ $O(t)$ $O(t)$ $O(t\cdot N^{2})$
Inference $O(t)$ $O(t)$ $O(t)$ $O(t)$ $O(t\cdot N^{2})$
4 Traditional Sequential ML Methods
With the increasing demand for higher data transmission rates and the limitations of traditional prediction methods reaching their practical limits in terms of accuracy, the need for algorithms with high precision, reliability, and low complexity has become urgent. In this section, we introduce a solution in DL to address this challenge. In summary, we begins with an overview of relevant research studies, subsequently providing a chronological exploration of key sequential models, namely, recurrent neural network (RNN), long short-term memory (LSTM), gated recurrent unit (GRU), and convolutional neural network (CNN). These architectures, which serve as fundamental components in DL models, are systematically introduced, along with concise examples that clarify their design principles, application methodologies, and complexity analysis.
In 2018, Karanov et al. BKaranov (2018) introduced an end-to-end deep neural network system for optical communications, encompassing the entire chain of a transmitter, receiver, and channel model. This pioneering work demonstrated the feasibility of transceiver optimization in an end-to-end manner. Owing to the intrinsic sequential structure of communication systems, sequential models, including LSTM networks GravesA (2012), RNNs PascanuR (2013), and GRUs JingL (2017) have been extensively employed since their inception. They continue to serve as crucial baseline models and foundational elements for developing more efficient algorithms.
RNN: It is originally designed for machine translation in natural language processing. This model is based on the Markov assumption about the hidden state and output sequence: namely, the output sequence depends only on the current potential state $h_{t}$. The potential state depends on the previous moment’s latent $h_{t-1}$ and input variables $x_{t-1}$ rather than on the historical data $\{x_{(t-1,\dots,0)},h_{(t-1,\dots,0)}\}$. Renowned for their adaptability in handling variable-length sequences and preserving state information across elements, RNNs find valuable applications in diverse communication fields XieY (2022). In recent work, RNN has shown promising results in equalization than the previous joint methods based on Volterra and Viterbi equalizers in two-dimensional eight-level pulse amplitude modulation (2D-PAM8) links X. Qin (2023). Despite the promising performance, it suffers from exploding gradient issues caused by the direct gradient flow of multiple layers Pascanu (2012). In other words, when the largest singular value of the gradient matrix with respect to parameters is larger than $1$ in the backpropagation process, the numerical instability will result in the training difficulty IEEE (1994).
LSTM: In contrast, LSTM addresses this issue by introducing input and forget gates to determine the gradient flow jointly. These additional states facilitate the learning of a weight matrix that preserves useful information in the hidden state Pascanu (2012). Here, we briefly recapitulate the architecture of LSTM; for more details, please refer to the original publication Sepp Hochreiter (1997). The input gate decides the information from the current input stored in the remote state, the forget gate regulates the retention or discarding of data from the prior hidden state, and the output gate defines the information presented as the final prediction. This gating mechanism equips LSTMs with the ability to effectively preserve pertinent details from lengthy sequences while discarding extraneous information. It is noteworthy, however, that LSTM does not guarantee a solution without vanishing/exploding gradients. Still, it does provide a more accessible way for the model to learn long-distance dependencies. LSTM has been widely applied in short-reach communication, especially for tasks that require the ability to capture intricate long-term dependencies LingP (2022).
GRU: Similar to LSTM, GRU is widely utilized as a variant of RNNs. It simplifies four hidden states into two critical gates, namely the update gate and the reset gate DeyR (2017). The update gate determines the relevant information to retain from the prior hidden state and the current input, while the reset gate dictates the data to discard. The resultant hidden state synthesizes the information preserved by the update gate and the current input. From an application perspective, GRUs find utility in scenarios where temporal dependencies and adjacent symbols within sequences play a crucial role. For instance, in short-reach communication, GRUs can alleviate signal distortions caused by chromatic dispersion and nonlinearities StavrosD (2021). For example, recent work has shown that the quality(Q)-factors can exceed the hard-decision forward error correction (HD-FEC) limit of 8.52 dB with the aid of bi-GRU nonlinear equalizer (8.52 dB estimated from $Q=20\log_{10}(\sqrt{2}\text{erfc}^{-1}(2\text{BER}))$)  LiuX (2021) .
CNN: Although CNN does not strictly belong to sequence models, it is widely applied in various domains due to its high parameter efficiency, weight-sharing mechanism, and plug-and-play characteristics NWu (2019). Convolutional operations use a convolutional kernel to scan the input signal with a fixed step on a specific dimension to capture the temporal features that are important for the downstream task. Typically, it is succeeded by a pooling layer and a nonlinear activation function, serving as redundancy reduction and nonlinear transformation, respectively. Essentially, features extracted by the convolutional kernel that closely align with the input signal are more likely to be preserved in the weight matrix. The relevant features crucial for downstream tasks will be enhanced by backpropagation processing. Recently, CNN has been utilized alongside belief propagation (BP) to improve decoding performance in an iterative manner Zhang (2023). Furthermore, a large number of test cases have shown that the combination of multiple layers of small convolutional kernels is often more efficient than using large kernels Szegedy (2014). Widely used and accepted foundational modules include inception and inception reduction, which extract temporal dependencies of different scales by employing concatenation of $1\times 1$, $3\times 3$, and $5\times 5$ convolutional kernels. In addition, a convolution kernel in the shape of $1\times 1$ serves a special function, thus reducing the channel dimension. It is generally used between two convolution layers or the output layer.
Summary: In this section, we have provided an introduction to the most common units of ML in optical communication systems. As of today, they are still directly used. In light of this, we summarize the aforementioned models’ complexity as follows. We neglect the choice of epochs and batch size but provide the complexity analysis for each batched sample.
Table 2: Complexity of DNN, GRU, LSTM, RNN and CNN. $t$ refers to the number of taps. $n_{s}$, $n_{o}$, $n_{h}$, and $d$ denotes input, output, hidden neuron, and depth of the DNN, respectively.
Models DNN GRU LSTM RNN CNN
Train $O(d-1)n^{2})$ $O((3n_{h}^{2}+6n_{h})n_{s})$ $O(4n_{h}^{2}+7n_{h})n_{s}$ $O(n_{h}^{2}n_{s})$ $O(n_{o})$
Inference $O(d-1)n^{2})$ $O((3n_{h}^{2}+6n_{h})n_{s})$ $O(4n_{h}^{2}+7n_{h})n_{s}$ $O(n_{h}^{2}n_{s})$ $O(n_{o})$
5 Advanced Sequential ML Methods
In Section 4, we introduced traditional sequence models, such as RNN, LSTM, GRU, and CNN. The key question we aim to answer in this section is how to effectively incorporate the unique characteristics of time series data into the modeling process and leverage the temporal convolution model to mitigate channel distortion. Different from other deep learning models to be introduced in the following sections, such as a transformer and Fourier-based Neural Networks, which necessitate meticulous parameter initialization and hyperparameter tuning when applied to new datasets, convolutional models exhibit superior generalization performance and are robust to parameter perturbation LIU (2022).
This section starts with channel modeling, encompassing four distinct noise models. We derive the characteristics and capabilities required for the algorithm based on these models. Subsequently, we provide a detailed exposition of the architectures and fundamental assumptions underlying three models: Frequency-Calibrated Sampled-Interaction Neural Network (FC-SCINet) CShao (2014)), DLinear Zeng (2022) and LightTS ZhangT (2022).
5.1 Channel Modelling:
The main limiting factor for the equalization task in a short-reach/PON system is ISI as a result of Chromatic Dispersion (CD), sampling error (Jitter), frequency shift (chirp), and other nonlinear effects BKaranov (2018).
Chromatic Dispersion in an optical communication system is caused by different phase velocities with respect to frequency. It fundamentally constitutes a linear transformation, and its mathematical representation involves a differential equation that considers spatial position and time, which can be presented as
$\frac{\partial A}{\partial z}=-j\frac{\beta_{2}}{2}\frac{\partial^{2}A}{\partial t^{2}}$ (1)
where $A$ denotes the amplitude of the complex signal, $t$ denotes time, $z$ is the spatial position along the fiber, where the pulse pattern propagates BKaranov (2018), and $\beta_{2}$ is the dispersion coefficient. Following the Fourier transformation, we have
$D(z,w)=\exp(j\frac{\beta_{2}}{2}\omega^{2}z)$ (2)
$w$ is the angular frequency. In the time domain, it is primarily manifested by significant attenuation in the high-frequency components and rapidly changing components.
Jitter is caused by fluctuations in sampling time. It presents itself as signal distortion, exhibiting a comparable impact to superimposed interference signals that adhere to the Gaussian distribution. Timing jitter can be described as
$\begin{split}y_{*}(t)&=y(t)\sum_{n=-\infty}^{+\infty}\delta(t-n*t_{A}-\tau)\\ &=y(n*t_{A}+\tau)\end{split}$ (3)
where $\tau$ is the timing sampling error, where the correctly sampled value is $y(n*t_{A})$. This sampling error can be quantitively estimated as follows:
$|y(n*t_{A}+\tau)-y(n*t_{A})|\leq M_{1}|\tau|$ (4)
where $M_{1}$ is the first moment of the band-limited spectrum of Fourier transformation of original signal $Y(f)$, can be simply written as:
$|\frac{\partial y(t)}{\partial t}|\leq\sum_{-f_{g}}^{f_{g}}|2\pi f||Y(f)|=M_{1}$ (5)
Jitter typically manifests as high-frequency fluctuations in the amplitude of an individual signal, and this elevated-frequency perturbation constitutes a highly effective form of attack on neural networks dependent on low-frequency signals.
In conclusion, the error can be estimated as $|y(t_{n})-y(n\tau_{a})|\leq M_{1}|t_{n}-n\tau_{a}|=M_{1}|\tau_{n}|$. The error $|e_{n}|$ is bounded by $M_{1}\cdot|\tau_{n}|$ for a given $n$. The value of $e_{n}$ depends solely on $\tau_{n}$. Assuming that the timing error $\tau_{n}$ follows a statistical nature with $E\{\tau_{n}\}=0$ and $E\{\tau_{n}^{2}\}=\sigma_{\tau}^{2}$, it follows that the amplitude errors $e_{n}$ are statistically independent. Consequently, the error variance is then given by $E\{e_{n}^{2}\}\leq M_{1}^{2}\sigma_{\tau}^{2}$. For more details, please refer to Kiencke (2008).
Chirp is a signal whose frequency varies with time. Mathematically, it can be described as follows,
$s(t)=a(t)\cdot\exp[j(\omega_{0}\cdot t+\theta(t))]$ (6)
The frequency spectrum of this waveform is obtained as
[…]
Simplifying further:
[…]
In summary, all types of effects encountered in equalization issues, except for jitter, involve concurrent alterations in both the time and frequency domains. It is noteworthy that such changes are not statistically independent. Consequently, no single effect can be eliminated through straightforward nonlinear operations in a single domain.
Figure 1: Sampling modules in temporal convolution networks modified from CShao (2014); ZhangT (2022)
5.2 Temporal Convolution Neural Network
Deep learning-based equalizers fundamentally capture intricate, domain-specific nonlinear disturbances by employing linear transformations and nonlinear activation functions. Within the temporal convolutional neural network, the core modules comprise interval, continuous, and interaction sampling modules, alongside convolutional neurons and linear layers, as depicted in Fig. 1. Each of these modules offers practical flexibility for hardware implementation, due to their computational efficiency. Subsequently, we delve into three of the most efficient convolution-based sampling networks.
FC-SCINet: This novel approach introduces an improved series decomposition technique as a spectrum correction module. In conjunction with the interaction sampling module, it has proven to be a robust tool for mitigating CD and addressing various real-world channel effects CShao (2014).
Decomp: In the case of FC-SCINet, it utilizes a moving averaging filter with kernel size $w_{1}$ to extract low-frequency signals from the input. Additionally, high-frequency signals are obtained by calculating the residuals between the original and low-frequency signals. The final output signal is generated through a weighted linear combination of these two components, which is $\hat{x}$ defined as Eq.(9).
$\hat{x}=W_{s}^{T}x_{s}+W_{f}^{T}x_{f}$ (9)
The complexity of this module is $O(k)$, where $k$ is the size of the kernel and is independent of the input sequence length.
However, as demonstrated in the empirical study in CShao (2014), the performance of FC-SCINet in mitigating CD remains strong. Moreover, the plug-and-play nature, low complexity, and interpretability of FC-SCINet make it highly flexible for seamless integration with various other algorithms. The DLinear architecture is another impressively low-complexity yet high-performance design.
SCIBlock: The SCIBlock, a crucial component of FC-SCINet, exhibits remarkable capabilities as it iteratively decomposes a signal into sub-sequences at various scales while incorporating nonlinear transformations between adjacent layers. In contrast, the decomp block is restricted to enhancing fixed signal components and is limited to a single scale. From a mathematical perspective, the SCIBlock applies a hierarchical structure by systematically downsampling the input sequence into even-positioned and odd-positioned samples, denoted as $x_{even}$ and $x_{odd}$. Following the convolutional layer, the sub-sequences in adjacent layers are iteratively multiplied together utilizing exponential and multiplication operations, as shown in Eq. (10) and (11).
$x_{even}^{s}=x_{even}\odot\exp(\psi(x_{odd})),\quad x_{odd}^{s}=x_{odd}\odot\exp(\phi(x_{even}))$ (10)
$x_{odd}^{\prime}=x^{s}_{even}+\exp(\eta(x^{s}_{odd}))\quad x_{even}^{\prime}=x^{s}_{odd}-\exp(\rho(x^{s}_{even}))$ (11)
Here, $\odot$ represents an element-wise product, and $\psi$, $\phi$, $\eta$, and $\rho$ are independent 1D convolutional layers. The intermediate outputs can be presented as $x_{even}^{s}$, $x_{odd}^{s}$, $x_{even}^{\prime}$, and $x_{odd}^{\prime}$. Upon completion of the processing, the resulting sub-sequences are then reassembled and aligned back to their original positions within the original signal. Ultimately, all the sub-series are concatenated based on their original index in the raw sequence, as illustrated in Fig. 1. To sum up, FC-SCINet is a framework capable of efficiently learning local-dependent patterns. Its distinctive feature lies in performing interactive learning on sub-sequences with odd-even positions after odd-even sampling, allowing for a larger receptive field under the premise of using the same convolutional kernel.
DLinear: As previously mentioned in FC-SCINet, while the concatenation of the decomp module may not offer optimal equalization, it has exhibited superior performance in real-world datasets. Therefore, we will provide a brief introduction to this module. It first decomposes a raw data input into a low $\mathbf{x_{s}}$ and high frequency $\mathbf{x_{f}}$ signal. $\mathbf{x_{s}}$ is extracted by a moving average kernel. It is equivalent to filtering the signal using a $sinc$ function in the frequency domain. These two components are added in a linear combination form, expressed by $W_{s}$, $W_{f}$. The operation above is presented in Eq.(12).
$\mathbf{x_{s}}=\text{AvgPool}(\mathbf{x})\quad\mathbf{x_{f}}=\mathbf{x}-\mathbf{x_{s}}\quad\mathbf{x}^{\prime}=W_{s}\mathbf{x}+W_{f}\mathbf{x_{f}}$ (12)
By iterative decomposition with different kernel sizes, DLinear can be extended to a deeper network. The complexity is $O(kn_{s})$, where $k$ is the number of the layer, $n_{s}$ is the length of the model input. To simplify the complexity, the weight matrix $W$ could be replaced by the convolutional kernel.
LightTS: Both FC-SCINet and DLinear utilize only convolution and different sampling modules to capture the local and global dependencies. The LightTS architecture, detailed in ZhangT (2022), employs a multi-layer perceptron (MLP) structure to enhance predictive abilities.
Sampling: In contrast to SCIBlock, which samples the raw sequence using odd and even indices, LightTS introduces two generic sampling strategies: Interval Sampling and Continuous Sampling, as shown in Fig. 1. Interval sampling partitions time series data into non-overlapping sub-sequences based on fixed time intervals, as shown in Fig 1. This approach helps identify periodic patterns or regularities within the data while minimizing information loss. On the other hand, continuous sampling divides sequences into continuous sub-sequences, extracting data points continuously throughout the time series and preserving temporal continuity. This sampling method enables the capture of patterns within the period, ensuring a more comprehensive representation of the underlying dynamics. The subsequent section presents an MLP-based architecture to extract useful features from both the down-sampled sub-sequences and continuously sampled sub-sequences.
IEBlock: The IEBlock serves as the central module in LightTS, designed to effectively process the 2D matrix resulting from continuous sampling and interval sampling. This block comprises three essential components: 1) temporal projection, which identifies temporal features following continuous sampling; 2) channel projection, which captures inter-channel information following interval sampling; and 3) the exchange block, which integrates the information from the outputs mentioned above, facilitating information fusion, all of them utilize MLP as the nonlinear behavior learning module. The design of LightTS is notably concise, employing only two sampling modules and an MLP. On certain datasets, it surpasses the performance of FC-SCINet Zeng (2022).
The above models require structural adaption, additional preprocessing etc. while applied in the practise. One of the recent work which has been successful in PONs applications includes the FC-SCINet CShao (2014). Different from LSTM, which offers the advantage of ensuring information flows strictly from past to future, temporal CNN goes beyond this by modeling the global dependency between input and output, while also leveraging stacked causal convolution layers. Additionally, the FC-SCINet introduces interaction modeling, enabling the explicit capture of interactions between elements within a sequence, making it a more advanced alternative to LSTM. In addition to these benefits, convolutional CNNs and SCInet offer several advantages over LSTM:
  • Convolutional CNNs can identify patterns regardless of their position within the input sequence. This property makes them well-suited for tasks where the position or timing of relevant features is not fixed, providing greater flexibility compared to LSTM.
  • Convolutional CNNs excel at capturing local patterns and extracting relevant features from the input sequence. This ability is particularly useful for tasks that require identifying and recognizing specific patterns or motifs within the data.
  • Convolutional CNNs and SCInet architectures typically have fewer parameters compared to LSTM models. This reduced parameter count can make training and inference more efficient, especially when working with limited computational resources or when dealing with large datasets.
Recent Progress: Convolutional networks maintain an indispensable role in current time series prediction research and practical applications due to their high parameter efficiency, model stability, and robust theoretical foundation (Multiscale Decomposition). The complexity of these networks primarily relies on the number of layers and the size of the convolutional kernels. Presently, advanced designs like dilated convolution and inception are frequently integrated with other modules to construct sophisticated deep learning models, yet they are seldom employed independently. Nevertheless, temporal convolutional neural networks retain an indispensable edge in terms of the Performance-vs-Complexity ratio and prove straightforward to implement in hardware.
5.3 Transformer-based Network
The transformer framework emerged as the predominant approach for sequential data analysis after 2017. This shift in dominance is primarily due to the transformer’s remarkable predictive accuracy demonstrated in empirical studies, as well as its exceptional ability to capture intricate dependencies within long sequences VaswaniA (2017). While the transformer was initially designed for natural language processing tasks, it remains the most extensively utilized sequence model and serves as the fundamental framework for numerous models WuH (2022). However, in this section, we will comprehensively elucidate the design rationale behind each module. The overview of the time series transformer is illustrated in Fig. 2.
Figure 2: Transformer Architecture
Encoder-Decoder: The objective of an encoder-decoder is to discover a reduced-dimensional representation of the input at an intermediate layer and subsequently reconstruct the output with minimal error. By doing so, the encoder-decoder acquires the ability to compress and reconstruct the input data nonlinearly. The transformer model’s encoder component consists of N stacked sub-encoders. Each sub-encoder encompasses a self-attention layer, a feed-forward layer, a residual connection, and layer normalization. The initial sub-encoder processes the input embeddings vector, while subsequent sub-encoders take the output from the previous sub-encoder as their input.
Attention: The Scaled Dot-Product Attention mechanism is the key component aiming to aggregate information across different parts of the input sequence. Each input vector is transformed into three distinct vectors: Queries ($Q$), Keys ($K$), and Values ($V$). The process involves calculating the dot products of the queries with all keys, scaling them by the square root of the dimension dk, and applying a softmax function to obtain weights on the values VaswaniA (2017).
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (13)
The resulting matrix of outputs is obtained through a weighted sum of the values, where the weights are determined by the softmax-processed dot products of queries and keys. This attention module allows the model to focus on relevant parts of the input sequence, capturing local dependencies during the training process. A residual connection is applied around the two sub-layers, followed by layer normalization to maintain the information flow.
Transformers employ multiple attention heads to implement multiple perspectives for different parts of the input sequence, allowing the model to simultaneously learn and consider multiple perspectives on the input data. Equation 14 and 15 represent the functionality of the multi-head attention step. $head_{i}$ represents the single attention head. The final result of the multi-head attention is concatenating all the attention heads.
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{o}$ (14)
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (15)
Although the vanilla transformer has demonstrated remarkable performance in dealing with time series data, the computational complexity of its original attention mechanism hampers its ability to process long sequences effectively. To overcome this limitation, researchers have developed various attention mechanism variants. An example of such is the Locality-Sensitive Hashing (LSH) Attention mechanism, which was introduced as part of the Reformer architecture KitaevN (2020). The LSH Attention mechanism utilizes specialized hash functions to transform queries and keys, thereby facilitating the categorization of similar items into shared hash buckets. Through sorting tokens based on their hash codes, items with similarities are grouped together, enabling the aggregation of relevant information. To enable parallel processing, the sorted sequence is divided into chunks. Subsequently, attention mechanisms are selectively applied to these chunks and their neighboring segments, allowing for focused examination of localized portions. By employing a systematic framework based on hash coding, the LSH Attention mechanism significantly improves computational efficiency compared to the vanilla transformer. It effectively addresses the challenge of processing long sequences by reducing computational complexity while preserving performance.
Decoder: The decoder consists of several stacked sub-decoders. The ground truth follows a process similar to that of the encoder, being transformed into Query $Q$, Key $K$, and Value $V$ representations. Attention weights are calculated by comparing the Queries $Q$ from the decoder with every value $V$ in the encoder. This process is repeated in parallel across N sub-decoders, resulting in a final attention matrix. The attention matrix then undergoes a softmax operation, yielding probabilities for each value. In addition to the two sub-layers in each encoder layer, the decoder introduces a third sub-layer, performing multi-head attention over the encoder stack’s output VaswaniA (2017).
Throughout the decoding process, the model is auto-regressive, taking previously generated outputs as additional input for developing the following output. Residual connections and layer normalization are applied around each sub-layer, and a specific self-attention sub-layer modification is made to avoid positions from attending to subsequent positions. Combined with the offset output embeddings, this ensures that predictions for a given position depend only on known outputs at positions earlier in the sequenceVaswaniA (2017).
Attention Variants: The objective of the attention layer is to identify inter-dependencies and connections among various input embeddings, allowing the model to evaluate the importance of each element in relation to others. It explicitly computes relationships between different elements in the sequence, which can be insightful and intuitive for understanding how information flows through the model. Except for the problem of computational complexity, the mechanism in vanilla transformer VaswaniA (2017) still needs to be improved in processing inter-dependencies and periodicity of signal data. A new type of encoder in the Autoformer WuH (2021) replaces the original encoder, which applies series decomposition and autocorrelation to detect inter-dependencies, combining representations within the sub-series. The series decomposition component divides the original signal into two distinct parts: the seasonal component, which captures short-term patterns, and the trend component, which captures long-term behavior. This partitioning allows for identifying and representing both the short-term and long-term characteristics present in the time series data. Additionally, the auto-correlation mechanism utilizes Fast Fourier transforms (FFT) to compute correlations between the time series and its delayed version, providing insights into how the series relates to its past values at different time lags. The combination of series decomposition and auto-correlation effectively captures and represents the underlying trend and seasonality within the time series data.
Recent Progress: Within this section, we provide a comprehensive overview of the vanilla transformer and its module design, particularly within the domain of time series and traffic prediction. Over the years, substantial improvements have been made to enhance the transformer for accurate time series prediction. Notably, advancements have been achieved in reducing computational complexity while improving the effectiveness of the attention mechanism logsparse (2019) informer (2021). However, recent research has introduced compact models based on multi-scale transformation scaleformer (2023), which unexpectedly surpass the performance of existing meticulously designed models across various benchmarks. This development has sparked a renewed and significant debate on the fundamental structure of sequence models. In the subsequent sections, we will summarize and delve deeper into this particular model, shedding light on its intricacies and implications. For the latest work, please refer to Table 3.
5.4 Fourier Convolution Neural Network
In the previous section, we introduced models based on convolutional kernels and subsampling as fundamental modules. Their core principle involves decomposing signals into different scales in the time domain and subsequently applying nonlinear transformations to learn salient features. However, for the majority of real-world signals, transforming them into the Fourier domain is often more efficient. This efficiency is attributed to the following factors: 1) The majority of real-world signals are bandpass or lowpass, and in the Fourier domain, their dynamic range decreases from $n$ to $exp(-n)$; 2) Fourier transformation is a bijective transformation, ensuring energy conservation and controllable error in both forward and inverse transformations; 3) The computational complexity of existing fast Fourier transform algorithms, after improvements, is $O(nlog(n))$, making it convenient for hardware implementation.
In this section, we will introduce TimesNet WuH (2022), while TimesNet utilizes a frequency-attention mechanism and FreTS YiK (2023) explicitly performs non-linear transformations in the frequency domain. For extensive models please refer to Table 3. The fundamental concept of TimesNet involves transforming the initial signal into $k$ distinct 2D tensors instead of directly processing the original sequence. This approach empowers the model to effectively capture both intraperiodic and interperiodic variations within these fixed windows. A variant based this model has achieved ShaoFConv (2024)
TimesNet: An attention mechanism based on spectral amplitude is employed to determine the significance of signal segments at various frequencies. Simultaneously, across different temporal resolution scales, a shared convolution module is utilized to reconstruct nonlinear distortions introduced by the channel. It does not explicitly perform nonlinear transformations in the frequency domain; instead, it combines reconstructed signals at different window lengths through a linear combination. The FConvNet primarily comprises four key blocks: Component Detection, Alignment, ConvNet, and Reconstruction.
Component Detection: Identifying the most crucial k frequencies is based on the amplitude of Fourier coefficients. Subsequently, utilizing the selected components within the k frequency range, the signal undergoes sampling using a continuous sampling method, and these subseries are aligned into a two dimensional tensor.
Alignment: The aligned subseries are then fed into a convolution-based module, specifically Inception, to alleviate distortion caused by channel effects. Fourier coefficients pass through a softmax function to generate attention weights, which are multiplied by the output of each convolution module to produce the final output.
Fourier Attention: Fourier transformation, being a global operation, induces periodic oscillations throughout the entire signal when there is any alteration in amplitude. Substantial variations in the amplitude of the primary components result in significant fluctuations. Leveraging this characteristic, TimesNet utilizes the Fourier spectrum of the nonlinearly transformed signal to determine the attention value for each component.
Reconstruction: Finally, employing a residual form, we obtain the reconstructed individual sub-components multiplied by their respective attention values, denoted as $y^{\prime}$, and add them to the input signal $x$ to yield $y$.
FreTS: As the local characteristics vary, any time-domain-based processing is constrained by the limitations imposed by information bottlenecks. In comparison, FreTS explicitly utilizes frequency-domain features to directly mitigate distortion without directly manipulating the time-domain in its model architecture. FreTS is an MLP-based network that is able to effectively learn patterns of time series data in the frequency domain. As presented in YiK (2023), FreTS consists of two learners: Frequency Channel Learner and Frequency Temporal Learner. In fact, in the equalization problem, we do not have a so-called channel dimension but rather a stack of independent experiments. Thus, we only introduce a frequency domain MLP.
Frequency MLP: The frequency temporal learner aims to capture temporal patterns in the frequency domain. Specifically, for a complex number input $\mathcal{H}\in\mathbb{C}^{m\times d}$, the MLP aims to optimize the weight matrix $\mathcal{W}\in\mathbb{C}^{d\times d}$ and bias $\mathcal{B}\in\mathbb{C}^{d}$ so that final output $\mathcal{Y}\in\mathbb{C}^{m\times d}$ could approximately reconstruct the ground truth.
$\mathcal{Y}_{\ell}$ $=\sigma(\mathcal{Y}_{\ell-1}\mathcal{W}_{\ell}+\mathcal{B}_{\ell})$ (16)
$\mathcal{Y}_{0}$ $=\mathcal{H}$ (17)
The MLP in the frequency domain implies the global convolutions in the time domain. The detailed proof can be found in YiK (2023). An increasing work has demonstrated the feasibility of deep learning models in the frequency domain. Simultaneously, the corresponding computational complexity of frequency-domain processing has decreased from $O(n)$ to $O(nlogn)$ due to the reduction in the signal’s dynamic range. However, the advantages and disadvantages of networks in both the frequency and time domains remain inadequately explored. Due to space limitations, we offer a more detailed categorization, along with corresponding references and keywords, in Table 3 for researchers with specific interests.
Table 3: Benchmark Models
Models Efficient Techniques Literature
Transformer Attention Sparsity inductive bias logsparse (2019) LogTrans leverages convolutional self-attention for improved accuracy with $O(L(\log L)^{2})$ lower memory costs.
Low-rank property informer (2021) Informer selects dominant queries based on queries and key similarities.
Learned Rotate attention (LRA) Quatformer (2022) Quatformer introduces learnable period and phase information to depict intricate periodical patterns.
Hierarchical pyramidal attention Pyraformer (2022) Pyraformer proposed one hierarchial attention mechanism with a binary tree following the path with linear time and memory complexity
Frequency Attention FEDformer (2022) FEDformer: proposed the attention operation with Fourier transform and wavelet transform.
Correlation Attention WuH (2021)Autoformer: the Auto-Correlation mechanism to capture sub-series similarity based on auto-correlation and seires decomposition
Cross-dimension dependency zhang (2023) Crossformer utilizes multiple attention matrices to capture cross-dimension dependency
Architecture Triangular patch attention Triformer (2022) Triformer features a triangular, variable-specific patch attention with a lightweight and linear complexity
Multi-scale framework scaleformer (2023)Scaleformer iteratively refines a forecasted time series at multiple scales with shared weights
Positional Encoding Vallina Position VaswaniA (2017) $\cos$ and $\sin$ functions with a sampling rate-relevant period.
Relative Positional Encoding Zerveas (2021) Introduces an embedding layer that learns embedding vectors for each position index.
Model-based learned logsparse (2019) LogSparse utilize one LSTM to learn relative position between series tokens
Fourier-NN Time Domain Series Decomposition Zeng (2022) DLinear performs one linear series decomposition with multiple layers
Frequency Attention WuH (2022) TimesNet proposes the attention mechanism related to the amplitude of the signal
Frequency Domain Frequency MLP YiK (2023) FreqMLP performs MLP in frequency domain by leveraging the global view and energy compaction characteristic
TConv-NN Sampling Continous ZhangT (2022) and WuH (2022) both utilize continous sampling to split original signal into windowed subseries similar to short time transformation
Interval ZhangT (2022) Interval sampling with fixed step to extract periodic feature
Even-Odd/Multiscale CShao (2014) proposes one iterative multiscale framework where even and odd series are interacted between layers
Frequency Continous Zeng (2022) leverages series decomposition module in a iterative manner to decompose signal in frequency domain with $sinc$ function.
Negative sampling TCN (2019) custom loss function is employed in an unsupervised manner, wherein distant or non-stationary subseries maximize the loss, while similar subseries minimize the loss.
Feature Module MLP ZhangT (2022) applies an MLP-based structure to both interval sampling and continuous sampling for extracting trend and detail information.
Dilated convolutions GraphWaveNet (2019) leverages stacked dilated casual convolutions to handle spatial-temporal graph data with long-range temporal sequences
6 Model Compression
In recent years, the proliferation of large-scale ML models has significantly advanced state-of-the-art technology across various domains, ranging from natural language processing to computer vision. The surge in model complexity, often characterized by intricate architectures and many parameters, has driven the need for efficient hardware implementations to harness their full potential. The advent of single graphics processing units (GPUs) as a critical computational resource has been pivotal in this landscape, offering a parallelized architecture suitable for accelerating the training and inference processes LewJ (2019). The significance of deploying large machine-learning models on a single GPU lies in optimizing computational efficiency and reducing latency. Single GPU implementations facilitate parallel processing, enabling the simultaneous execution of multiple tasks and handling extensive model parameters. This enhances the speed of model training and expedites real-time inference, a critical requirement in applications such as autonomous systems and edge computing. However, the hardware implementation of large machine-learning models on a single GPU is challenging. The complexity of these models often exceeds the computational capacity and memory constraints of a single GPU, necessitating innovative solutions for efficient utilization MarculescuD (2018). Techniques such as model pruning, quantization, vector quantization, and knowledge distillation have emerged as strategies to mitigate these challenges, ensuring that even formidable models can be accommodated within the limitations of a single GPU without compromising performance. The authors examine how to use a single GPU effectively for implementing large machine learning models. They discuss methods that balance complexity and computational efficiency to maximize hardware utilizationMarculescuD (2018).
However, conducting a comprehensive performance-versus-complexity analysis is required to evaluate the suitability of various ANNs in short-reach optical communication systems. Deep learning models, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), and long short-term memory networks (LSTMs), find applications in critical tasks such as equalization, fault detection, subcarrier allocation, nonlinearity compensation, and bandwidth request and allocation. The complexity of these models is a significant factor affecting their feasibility and use in practice. For instance, CNNs, renowned for their effectiveness in image processing, may introduce convolutional and pooling layers, increasing model complexity. Similarly, RNNs and LSTMs, designed for sequential data, introduce recurrent connections that enhance their ability to capture temporal dependencies and contribute to increased complexity Freire (2021). Analyzing the intricacies of NN architectures, including their depth, the number of parameters, and computational demands, is crucial for understanding the trade-offs between performance and complexity. Deep learning models often exhibit enhanced capabilities in capturing intricate patterns and relationships in optical communication data. Still, their complexity may pose challenges regarding training time, computational resources, and practical implementation Asghar (2019). A thorough examination of these complexities is essential for identifying optimal models, such as choosing between a CNN for image-based tasks or an LSTM for sequential data, that balance high performance and manageable complexity, facilitating their efficient integration into short-reach optical communication systems Freire (2021). Four prominent types, namely feed-forward neural networks (FFNN), radial basis function neural networks (RBF-NN), auto-regressive recurrent neural networks (AR-RNN), and layer-recurrent neural networks (L-RNN), offer distinct trade-offs in complexity and performance. Numerical evaluations reveal that, for a fixed bit error ratio (BER) threshold, AR-RNN-based equalizers exhibit the lowest computational complexity. Among nonlinear NN-based equalizers with equivalent numbers of inputs and hidden neurons, FFNN-based equalizers generally boast the lowest computational complexity. In contrast, AR-RNN-based equalizers demonstrate superior BER performance Xu (2019).
Distillation model: Knowledge distillation, a model compression technique, transfers knowledge from intricate, large-scale models or ensembles to more compact, feasible models suitable for real-world constraints. Pioneered by Bucilua and collaborators in 2006 Buciluǎ (2006), knowledge distillation primarily operates on neural network architectures characterized by multifaceted structures comprising multiple layers and parameters. The surge of deep learning over the past decade, notably successful in domains like speech and image recognition and natural language processing Gou (2020), has propelled knowledge distillation as a pivotal methodology for practical applications. The demand for deploying extensive deep neural network models becomes particularly challenging for edge devices limited by memory and computational resources. In response, an innovative model compression method Gou (2020) emerged, facilitating the seamless transfer of knowledge from expansive models to train more resource-efficient counterparts without substantial performance degradation. This transformable process, where a smaller model acquires knowledge from a larger model, was systematized into the formal "Knowledge Distillation" framework by Hinton and colleagues Hinton (2015). This framework has become integral for addressing the computational constraints of edge devices while preserving the essential knowledge distilled from sophisticated, larger-scale models.
Optimizing deep learning models through knowledge distillation holds substantial promise for future advancements in short-reach optical communication systems. Notably, recurrent neural networks (RNNs) have demonstrated unparalleled efficacy in equalizing nonlinear impairments Freire (2021) StavrosD (2021). However, the inherent feedback loop in RNN structures poses a formidable challenge to parallelization, hindering their implementation in low-complexity hardware tailored for high-speed processing in optical networks AndreXianMingC (2017). To overcome this limitation, a prospective avenue involves harnessing knowledge distillation as a strategic solution for enabling parallelization in RNNs Freire (2021) SasipimS (2023). This application of knowledge distillation is poised to revolutionize the implementation of RNNs, ensuring compatibility with low-complexity hardware and meeting the stringent processing requirements of high-speed optical networks RobertR (2021).
Looking beyond RNNs, the future trajectory of knowledge distillation extends its influence across various machine-learning models crucial to optical communications. Examples include CNNs, LSTMs, FFNN, RBF-NN, AR-RNN, and L-RNN SasipimS (2023). These diverse models present distinct challenges in complexity, adaptability, and real-time processing. For instance, consider the scenario of implementing knowledge distillation in CNNs designed for optical signal processing tasks. The distillation process can facilitate knowledge transfer from a more complex teacher model to a simplified student model, optimizing the trade-off between accuracy and computational efficiency. Similarly, in LSTMs utilized for sequence processing in optical communication systems, knowledge distillation can streamline the model complexity without compromising its ability to capture intricate temporal dependencies SasipimS (2023).
Another field that holds promise in short-reach optical communication systems is knowledge distillation techniques that address challenges with limited time-series data. As "big data" impacts various fields, the scarcity of target events or high acquisition costs hinders machine learning tasks in specific scenarios. The proposed method, rooted in generalized distillation, utilizes "privileged information" from middle-time data during training. This novel approach enhances long-term predictions for small time-series datasets. Applied to optical communications, it offers a solution to data constraints, demonstrating efficacy on synthetic and real-world data. This distillation model presents a pathway for more robust predictive modeling in optical communication systems ShogoH (2019).
Vector quantization: Vector quantization (VQ) is a model compression technique to mitigate the complexity of large-scale machine learning models. The essence of VQ lies in its ability to represent intricate data with a compact set of prototype vectors, thus significantly reducing the computational load during inference. This technique finds applications where limited resources require a balance between model efficiency and performance. The working principle of VQ involves partitioning the input space into regions, each associated with a representative prototype vector. During the encoding phase, input vectors are assigned to the nearest prototype, effectively quantizing the data. In contrast, the given prototype vectors are employed to reconstruct the original data during the decoding or reconstruction phase.
The efficiency of VQ hinges on the strategic selection and adaptation of prototype vectors, optimizing their representation to encapsulate the essential information within the dataset RobertG (1984). By clustering and quantizing input vectors into a representative codebook, VQ enables encoding information in a more compact form. This is particularly beneficial in scenarios with limited data availability or high computational demands. For instance, consider the application of VQ in knowledge distillation for recurrent neural networks (RNNs), which face challenges in parallelization due to their feedback loop structure. Implementing VQ in the distillation process can help mitigate the parallelization issue by representing the essential information with a reduced set of prototype vectors XinW (2019) KashifR (2022). This compression not only aids in overcoming hardware complexity but also contributes to faster processing in high-speed optical networks.
VQ typically involves an iterative process that refines the prototype vectors to enhance their representative capacity. This is commonly achieved through algorithms like the k-means clustering algorithm, where prototypes are adjusted to minimize the overall distortion or reconstruction error between the original data and the quantized representation. The iterative refinement of prototype vectors allows vector quantization to adapt to the inherent patterns and structures within the data, optimizing its compression capabilities while preserving critical information for training tasks RobertG (1984).
Furthermore, VQ can be crucial in optimizing other deep learning models, such as CNNs or LSTMs, by efficiently capturing essential features with a minimal set of representative vectors SherjilO (2021). The future exploration of VQ in conjunction with these models offers a promising avenue for improving the performance and scalability of machine learning applications in short-reach optical communication systems.
7 Conclusion
In this survey, we have undertaken a comprehensive examination of powerful machine learning models that exhibit the potential to achieve robust equalization in cost-sensitive short-reach systems, with a particular focus on PONs. Our objective has been to explore these models’ capacity to operate efficiently and deliver effective computational performance. For the first time, we have classified the current models into three distinct types and conducted an extensive analysis of their core concepts, highlighting their differences, similarities, and the underlying insights they provide. Additionally, we have presented a simplified complexity analysis considering various input sizes. In the final stages of our survey, we have also investigated the potential of machine learning solutions in addressing the challenges associated with hardware implementation and complexity. We firmly believe that this survey will serve as a valuable resource, inspiring future research endeavors to develop efficient models explicitly tailored for short-reach and PON systems.
{adjustwidth}
[…]